6x-2y= -4 What would I need to do to find the answer


Answer by 121086(1)



6x-2y = -4
-6x+2y = -4 - 6x First start by adding -6x to both sides leaving the y
variable
+2y = -4 - 6x Then divide by 2 so the coefficient in front of y equals 1.
(1)y = -2 - 3x 2 divided by 2 equals 1, 2 divided by -4 leaves -2,
& 2 divided by -6x leaves -3x.
y = -2 -3x